Hardware Consultants purchased a building for $ 510,000 and depreciated it on a​ straight-line basis over a 30​-year period. The estimated residual value is $ 90,000. After using the building for 15​ years, Budget realized that wear and tear on the building would wear it out before 30 years and that the estimated residual value should be $ 80,000. Starting with the 16th​ year, Budget began depreciating the building over a revised total life of 25 years using the new residual value. Journalize depreciation expense on the building for years 15 and 16. ​(Record debits​ first, then credits. Select the explanation on the last line of the journal entry​ table.)
Answer #1

Question is asking only for the entry for expense, so closing entries are not passed below,

Let us begin the answer by calculating depreciation for both the asked years.

  1. 15th year, this is as per first method the company has followed, cost = 510,000, life = 30 years, residual value = 90,000, method = SLM. So depreciation = (510,000 - 90,000)/30 = 14,000.
  2. 16th year, here the depreciation amount will differ, because useful life and residual value will be changed, as cost we should consider WDV after 15 years, that is 510,000 - depreciation fir 15 years = 510,000 - 210,000 = 300,000. So depreciation = (300,000 - 80,000)/10* = 22,000. *Remaining life is 10 years.

Journal entries,

15th year,

Particulars

Debit Credit
Depreciation expense 14,000
Building 14,000

Being depreciation expense recorded for the year 20xx

16th year,

Particulars Debit Credit
Depreciation expense 22,000
Building 22,000

Being depreciation expense recorded for the year 20xx